Rivendell, a small open economy, manages to have a national savings equal to 1,001. Its investments are worth 500. If the equation for net export is given by

NX = 2500 - 200ε

what is the real exchange rate? Round to the nearest hundredth decimal place (2 decimal places).
Answer #1

Answer

In an open economy Saving-income equation(identity) is given by :

S = I + NX , where S = National saving = 1001, I = investment = 500 and NX = Net exports = 2500 - 200ε

=> 1001 = 500 + 2500 - 200ε

=> ε = 9.995 = 10(round to teo decimal places)

Hence, Real exchange rate = 10